Express in formality the concentration of 36.0 per cent HCl by weight heaving a density of 1.19 grams per ml


1
Expert's answer
2021-02-23T04:29:27-0500

CM=n/V

n=m/M

M (HCl) = 36.46 g/mol

Lets assume that we have 1 L of HCl solution.

density = m/V

m = density x V

m (HCl)solution = 1.19 x 1 = 1.19 kg

%= (100 x m(solute))/(m(solute)+m(solvent))

m(HCl) = (36.0 x 1190)/100 = 428.4 g

n(HCl) = 428.4/36.46 = 11.8 mol

CM(HCl) = 11.8/1 = 11.8 mol/L